Zephyr Farming Pty Ltd is considering the purchase of a wind turbine generator in order to generate electricity and to reduce the electricity costs for their offices, which are located in Toowoomba. Currently the business uses 60,000 kilowatt hours (kWh) per quarter (3 months) at an average cost of $0.30 per kwh, supplied by the local coal fired power station. The current required rate of return used to evaluate projects is 6%, with a required payback period of 3 years.
The Queensland government started a scheme to provide an incentive for business to use alternative sources of renewable power. The incentives are 5% immediate reimbursement of the purchase and installation costs. This reimbursement can be paid to the supplier providing and installing the equipment. Therefore, the net cash flow from Zephyr Farming Pty Ltd is the purchase cost plus the installation cost less the 5% incentive back from government to offset these costs.

Project details:
Cost of wind turbine generator                     $6,000
Cost to install turbine and generator (by supplier)             $450
Expected cash incentive back from government to offset cost of the panels paid immediately the wind turbine generator installed                   5% of total costs
Turbine expected (on average) generated kilowatt     300 hours per month
Generator’s expected life (in years)     15 years

Requirements
1. Calculate the total initial investment and the net annual savings from installing the wind generator. (8 marks)
2. Calculate the payback period of the wind generator. (4 marks)
3. Comment on whether this project meets the company’s Payback Period requirements and whether you should base your decision only in the payback period calculation.
(1 mark)
Answer #1

1. Cost of turbine generator = $6,000

Add: installation cost = $450

Less: Reimbursement from government = 5%*6,450 = $322.5

Total initial investment = $6,127.5

Calculation of net annual savings:

Generation of kilowatt hours per month = 300

Annual = 300*12 = 3,600

Cost per kWh = $0.30

Net annual savings = 3,600*0.30 = $1,080

2. Payback period is the time period in which the initial cost is recovered.

Payback period of generator = 6,127.5/1,080

= 5.67 years

3. The acceptable payback period for company is 3 years, while the payback period for generator is 5.67 years. Hence, the generator doesn't meet company's payback period requirements.

It is not right to base the decision on only payback period. The generator will provide benefits for 15 years and required return is 6% on investment.

Based on this information, the generator will provide benefits for a long term and hence is a good investment.